OpenStudy (anonymous):

Solve the following system using substitution. x+3y=8 2x-y=2

OpenStudy (istim):

Rearrange one of them to solve for one of the variables (x or y)

OpenStudy (istim):

Then, sub in the rearranged equation into the unmodified question.

OpenStudy (istim):

From there, solve for the unknown variable.

OpenStudy (istim):

I think our friend here will give you more specified information if you need it.

OpenStudy (anonymous):

x = 8-3y so; 2(8-3y) - y = 2 -----> 16-7y = 2 so , y = 2 x = 2

OpenStudy (anonymous):

Thank You So Much I UNDERSTAND it now.